OpenStudy (anonymous):

WILL REWARD MEDAL TO THE BEST ANSWER. A Golf ball is struck so that it leaves the head of the club with a speed of 50.0m/s at an angle of 30.0° to the ground. Determine how high and how far the golf ball moved at 1.00s, 2.00s, 3.00s, 4.00s and 5.00 seconds. Please give the horizontal and vertical displacement in metres :)

OpenStudy (shamim):

Ur golf ball is a projectile

OpenStudy (shamim):

Given v0=50m/s Theta not=30 degree

OpenStudy (shamim):

The formula for horizontal displacement is x=v0*cos theta not*t

OpenStudy (shamim):

For ur vertical displacement y=v0*sin theta not*t-(1/2)*g*t^2

OpenStudy (shamim):

If u need more explanation then i wanna give u an attachment